Calculate the value of the following bond in the secondary market.

Face Value= $20,000, 7% semi-annual bond with five years remaining until maturity.The current rate on new five-year issues of comparable quality is 8%.
Answer #1

Value of bond is equal to the present value of all future coupon payments and the redemption value

Semi-annual rate = 8%/2 = 4%

Number of semi-annual periods = 5*2 = 10

Hence. value of bond = 20,000*7%*1/2*PVAF(4%, 10 periods) + 20,000*PVF(4%, 10 periods)

= 700*8.1109 +20,000*0.6756

= $19,189.63